OpenStudy (roseblooms):

Alana raked leaves for 7 days in the fall. She raked 9 bags of leaves each day. Alan earned d dollars for each bag she raked. Which expression could be used to determine the total amount of money she earned for the 7 days? Question 22 options: 63/d 63d 9d + 7 (9 + 7)d

OpenStudy (shamim):

1 day=9 bags of leaves 7 days=7*9=? Bags

OpenStudy (shamim):

1 bag= d dollar 63 bag=? Dollar

OpenStudy (sharmagk83):

Total money earned is given by \[= 7\times9\times d = $ 63 d\]